# QuFeX: Quantum feature extraction module for hybrid quantum-classical deep neural networks

### v1

@@ I Introduction
Quantum computing has emerged as a computational paradigm with the potential to solve problems intractable for classical computers. The rise of the machine learning (ML) methodology has spurred interest in quantum machine learning (QML) as a potential means to improve classification, regression, and clustering tasks Abohashima et al. (2020); Schatzki et al. (2021).
@@
Early on in the development of QML, quantum neural networks (NNs) were proposed Cong et al. (2019); Henderson et al. (2019) as an extension of the concept of classical NNs by utilizing quantum circuits as layers of the network. The design of the quantum layers corresponds to different ML architectures, with notable examples including quantum convolutional NN (QCNN) Cong et al. (2019) and Quanvolutional NN (QuanNN) Henderson et al. (2019). While convolutional NNs (CNNs) – the cornerstone of modern NNs – have achieved significant success across a wide range of ML applications Li et al. (2022a); Younesi et al. (2024), their resource requirements, especially for training large models, increase drastically with the size and complexity of data Thompson et al. (2022); Hestness et al. (2017). Quantum NN extensions, such as QCNN and QuanNN, were proposed under the hypothesis that, since quantum computers can process information in fundamentally different ways, quantum NNs may potentially offer greater efficiency compared to their classical counterparts Schuld and Petruccione (2018); Dunjko and Briegel (2017).
@@
In parallel to developing quantum NN models, in recent years there has been a growing interest in creating a hybrid quantum-classical model where quantum and classical NNs work in tandem Zaman et al. (2024); Hafeez et al. (2024); Zhang et al. (2024); Arthur and Date (2022). This division of computing resources allows leveraging the strengths of both classical and quantum computing while managing the limitations of near-term quantum devices. Hybrid models can be particularly promising in domains such as image processing and generative modeling, where quantum circuits may offer advantages in high-dimensional feature spaces, while classical systems handle more routine tasks. Thus, the hybrid quantum-classical ML approach provides a feasible pathway toward enhancing ML tasks under the current quantum technological constraints.
@@
A typical hybrid model includes classical layers for pre-processing data, followed by quantum processing layers (circuits), and concludes with classical post-processing steps. Nevertheless, various strategies have been proposed for integrating quantum layers into classical deep NN architectures. For example, in image classification tasks, quantum layers are often positioned near the end of the hybrid network to generate the final output label Senokosov et al. (2024), while some approaches involve encoding data through quantum layers at the beginning of the network Li et al. (2022b). The design of the quantum layer itself varies widely as well, with notable examples including QCNN Cong et al. (2019) and QuanNN Henderson et al. (2019). These variations underscore the need for a comprehensive study of optimal quantum layer placement in hybrid networks, as well as a systematic approach to constructing effective quantum layers.
@@
In this work, we take a step forward in these two directions. We introduce a novel quantum learning architecture, which we term quantum feature extraction, or QuFeX for short, which combines and leverages techniques from QCNN Cong et al. (2019) and QuanNN Henderson et al. (2019). At a high-level it integrates the data analysis structure of QCNN, which was shown to excel in classification tasks Elhag et al. (2024); Farhi and Neven (2018); Hur et al. (2021), with the data pipeline of QuanNN which is well suited for local transformation of the input data Henderson et al. (2019). This combination allows us to design a quantum layer that can extract features from the data in a reduced dimensional representation of it. This in turn results in a reduced number of parallel quantum circuit evaluations compared to QuanNN. Thus, one of the main advantages of the QuFeX design is in its ability to be integrated in a hybrid fashion with classical layers, which makes it an excellent candidate for integration in deep classical NNs.
@@
As a proof of concept, we propose a novel hybrid quantum-classical model, a Qu-Net, were the QuFeX is integrated at the bottleneck of a classical U-Net model. U-Nets are a type of convolutional neural network architecture primarily designed for image segmentation tasks Ronneberger et al. (2015). They are widely used in applications such as medical imaging Ronneberger et al. (2015); Yuan and Cheng (2024), autonomous driving Iglovikov and Shvets (2018), and satellite image analysis Bastani et al. (2018), where precise pixel-level predictions are essential. Our findings thus provide numerical evidence that with a careful design of quantum learning circuits, and with a strategic placement of those within deep NNs we can achieve quantum-classical hybrid learning models that leverage the strengths of both computing paradigms. Our demonstrations represent one of the first applications of quantum-enhanced deep NNs to real-world image segmentation tasks.
@@
The remainder of the paper is organized as follows. In Section II, we first present a brief overview of CNNs and their quantum counterpart architectures, setting the stage for the introduction of the QuFeX module. In Section III, we introduce the Qu-Net architecture utilizing the QuFeX at the bottleneck of the network. We numerically test and analyze the performance of the Qu-Net , for various model sizes, compared to an all-classical U-Net baseline. Finally, we offer conclusions and outlook in Section IV.
@@ II.1 Convolutional neural networks and their quantum counterparts
Motivated by these powerful information processing components researchers have developed quantum counterparts of CNN. QCNN, first introduced in Cong et al. (2019), was designed to mimic the translational invariance property of CNN. Figure 2 provides an illustration of the QCNN architecture. In QCNNs, the (classical) convolutional layers are replaced with a set of parametrized unitary transformations ($U_{j}$ in the figure) and pooling layers are replaced with fixed unitary operations ($V_{j}$ in the figure) conditioned on the measurement results of few specific qubits. These blocks of unitary operations (circuits) are concatenated in a serial fashion, and as shown in Fig. 2, each unitary circuit is designed in a translational invariant way to maintain this CNN property. One of the advantages of this approach is that for an $N$ qubit circuit, this network has only $O(\log N)$ trainable parameters Cong et al. (2019). While the structure of each circuit layer is fixed, the parametrized unitary operations ($U_{j}$) themselves are trainable via an optimization strategy, e.g., gradient descent to minimize a particular loss function, and therefore are often thought of as the quantum-equivalent of CNN kernels.
While, operationally, the quantum convolutional circuits within QCNNs are considered efficient data processing modules that mimic the translational invariance of kernels in CNN, the data pipeline of QCNN generally does not possess some of the key features of CNNs. Specifically, since the convolutional and pooling layers in a typical QCNN circuit are ordered sequentially a circuit, the input to the circuit is often too large to be handled efficiently (this is known also as the input problem). On the other hand, classically manipulating the input to reduce it to a manageable size, may result in loss of information when when handling data with multiple (or non-local) features.
@@
To address these challenges, QuanNN Henderson et al. (2019) has been proposed, which leverage quantum circuits to enable the simultaneous extraction of local features across input data. QuanNNs act as feature extractors, analogous to convolutional layers in classical networks, and can be integrated seamlessly with classical NN architectures for hybrid quantum-classical processing. As illustrated in Fig. 3, in QuanNN each quantum circuit is treated as a kernel and generates a feature map when applied to an input by transforming spatially-local subsections of it. These filters correspond to specific circuit designs and can be configured to potentially learn different features of the input. The QuanNN architecture thus offers greater flexibility and is considered a closer representation of classical CNNs in its way of handling the flow of data. Nonetheless, by mimicking the data flow of classical CNN, QuanNN fall short in efficiency. Since QuanNNs extract local quantum features independently for each patch, similar to CNNs, to process the input data we must design ${\cal O}(kN)$ quantum experiments, with thousands of learning iteration per filter as reported in  Henderson et al. (2019), where $N$ is the size of the input and $k$ is the number of filters used.
Figure 4: Example of a QuFeX architecture. The QuFeX incorporates ideas and techniques from QCNN and QuanNN. In the data flow we use information taken from different feature maps (from two maps in this example) and pass it to a convolutional-inspired quantum circuit. Example of such circuit is illustrated on the top-right part of the figure where $U_{1}$ and $U_{2}$ depend on trainable parameters (denoted in the paper by $\vec{\theta}$), see Appendix for more details. Single-qubit measurements are then performed and the average of the corresponding Pauli-$Z$ operator is then passed on to produce an output feature map. In the QuFeX architecture we use different circuits (i.e., circuits on different qubits) to analyze different patches of the input data, in parallel. The circuits that run in parallel contain the same trainable set of parameters.
@@ II.2 QuFeX: Quantum feature extraction architecture
In this architecture, the QuFeX circuit is designed as a quantum filter that creates a feature map from its input data. To avoid limiting the application of the filter to a spatially local subsection of the input data, we design it to act on a mix of different input feature maps, as illustrated in Fig. 4. This enables us to generalize the quantum feature extraction to the important (and common) cases where features are not spatially local in the input data. The way we mix the input data can be considered a hyperparameter, similar to CNN. Notably, this mixing can be done in parallel on all feature maps input, and therefore, the QuFeX can handle $k$ input feature maps with $O(1)$ uses of the QuFeX circuit [QuFeX$(\vec{\theta})$ in the figure]. This is one of the main advantages of the proposed architecture, compared to existing ones. By using QuFeX circuit on a small number of qubits (say, 4 qubits) we are able to handle $k$ classical feature maps, where $k$ can be very large, each of size $N\times N$, to produce an output feature map of size $M\times M$ (typically $M$ is chosen to be smaller than $N$). The cost of producing an output feature map using QuFeX depends only the size of the QuFeX circuit (which is logarithmically in the number of qubits) and the output feature map size $M$. The exact cost may vary from one implementation to another, depending on the choice of hyperparameters such as how we mix different input feature maps and the way we slide the convolution window. In addition, more filters, and hence more output feature maps, can be created simply by adding multiple QuFeX layers, which can run in parallel and with each being fully customizable.
Within each QuFeX layer we process the data in a translationally invariant way, similar QCNNs, as depicted in the Fig. 4, with one critical change. Here, we are not discarding the control qubits in the pooling layer. In QCNN the pooling layers are implemented by discarding the measured (control) qubits, effectively reducing the dimension of the network. In the QuFeX architecture, similar to classical CNNs and QuanNN, the reduction of dimension is done directly through the data flow. Therefore, instead of discarding the control qubits, and by doing so discarding information within a filter, we keep the control qubits and treat their output as a part of the (or even separate) feature map.
@@
As a proof of concept, in the next section we propose a novel hybrid quantum-classical model, a Qu-Net, were the QuFeX is integrated within a classical U-Net model Ronneberger et al. (2015)— a CNN with a U-shape topology specifically developed for image segmentation. We demonstrate the potential power of the QuFeX though segmentation performance of the Qu-Net architecture. Our demonstrations represent one of the first applications of quantum-enhanced deep NNs to real-world image segmentation tasks.
III The Qu-Net architecture with application to image segmentation
@@ III.1 The U-Net architecture
Illustrated in Fig. 5(a), the U-Net architecture is a deep CNN with a U-shape topology specifically developed for image segmentation Ronneberger et al. (2015). It features an encoder-decoder structure with skip connections, which help preserve spatial information, and enable precise pixel-level predictions for more accurate segmentation results. The encoder progressively downsamples the input image, capturing essential features through convolutional and pooling layers. Subsequently, the decoder upsamples the compressed feature map to the original input size, reconstructing the segmentation map. At its bottleneck, i.e., at the most compressed layers, the U-Net transforms features into a latent representation that bridges the encoder and decoder. This design “encourages” the model to retain only the relevant information in the bottleneck layer, facilitating the construction of an effective segmentation map.
This design helps U-Net excel in segmenting objects with fine boundaries, making it highly effective, for example, in medical applications like tumor delineation, organ segmentation, and cell tracking Ronneberger et al. (2015); Yuan and Cheng (2024). Beyond medical imaging, U-Net is used in various fields requiring pixel-level classification, e.g., for satellite image analysis Bastani et al. (2018) and autonomous driving Iglovikov and Shvets (2018). Its ability to handle small datasets through data augmentation and its versatile performance in capturing detailed structural information has made U-Net a widely adopted tool in image segmentation across industries.
@@ III.2 The Hybrid Qu-Net architecture
The downscaling of the input all the way down to the bottleneck of the architecture, introduces limitations on the representability of the original image. In addition to QuFeX, inspired by the ResNet architecture He et al. (2015), we include (classical) residual connections from one part of the network to the other that bypass the QuFeX module, see Fig. 5(b). The residual connection can be expressed as $\mathbf{y}=\mathcal{Q}(\mathbf{x})+\mathbf{x}$ where $\mathcal{Q}(\mathbf{x})$ represents the output of the quantum layer applied to the input $\mathbf{x}$ and the addition of $\mathbf{x}$ serves as the identity mapping. We argue that residual connections, typically used to stabilize training and convergence in deep NNs, are very relevant in hybrid quantum-classical setups. Adding identity mapping over the quantum layer enables the model to propagate features learned by the classical layers forward and backward through the network without being disrupted by the quantum transformation. This configuration allows the network to select an optimal path: either through a classical-only route or a path incorporating quantum processing, thus leveraging the advantages of both. Our tests below show that introducing residual connections to the Qu-Net architecture provides better segmentation performance compared to baseline U-Net models where the QuFeX is not included.
III.3 Testing and evaluation
@@
We have tested and evaluated the performance of the proposed QuFeX module within the novel Qu-Net architecture using the FruitSeg30 Segmentation Dataset & Mask Annotations Shamrat et al. (2024) which consists of high-resolution images of a variety of fruits and their corresponding segmentation masks. The complete dataset includes 30 fruit classes, totaling 1969 images, each paired with precise segmentation masks. For this research, a curated subset of $751$ images was selected (along with their segmentation masks) and downscaled, using Bilinear interpolation, from the original $512\times 512$ pixels to $64\times 64$ pixels to align with the available computational resources [Intel Core i5-11300H CPU @ 3.10GHz (4 cores), 8 GB DDR4 RAM running Ubuntu 22.04 LTS]. The subset was selected to ensure diversity between fruit types while maintaining computational feasibility.
To provide a baseline for the performance of the Qu-Net architecture, a U-Net architecture with approximately the same number of trainable parameters was also trained and evaluated on the same dataset. The code for all the numerical experiments reported in this subsection are available in a public repository Jain (2024).
We tested three variations of the models, which we refer to as tiny, small, and medium, that consist of approximately 12,000, 26,000, and 40,000 trainable parameters, respectively. These variations in parameter count were achieved by adjusting the number of filters in the classical layers of the architectures. We refer the reader to the Appendix for more details about the specific U-Net and Qu-Net models design.
@@
Our models were trained using an Adam optimizer with a learning rate of $0.001$ and a binary cross-entropy loss function. Training was performed for 10 epochs with a batch size of 64. The Qu-Net architecture was executed on quantum simulator using PennyLane et al. (2018). An ablation study was conducted to evaluate the influence of the QuFeX layer’s design choices, see the appendix for more details. The performance of Qu-Net was analyzed with varying numbers of quantum filters and qubits per filter within the QuFeX module. Specifically, here we report the results obtained using a quantum filter with 8 qubits and 2 quantum filters with 4 qubits per filters. We denote those models by Qu-Net $8(1)$ and Qu-Net $4(2)$, respectively.
@@
The statistical summary of these metrics is visualized in Figs. 7 and 8, providing a comparative analysis of the models’ performance. Each boxplot depicts the distribution of the results (IoU) from the 10 independent runs for the tested models. The median value is marked, with the interquartile range (IQR) capturing the variability of performance across different dataset partitions. Outliers are displayed to identify cases where the model’s performance deviates significantly from the norm. The complete distribution of results is given in Fig. 10 in the appendix.

### v2

@@ I Introduction
Quantum computing has emerged as a computational paradigm with the potential to solve problems intractable for classical computers. The rise of the machine learning (ML) methodology has spurred interest in quantum machine learning (QML) as a potential means to improve classification, regression, and clustering tasks qml_survey_2020; entangled_dataset_2021.
@@
Early on in the development of QML, quantum neural networks (NNs) were proposed iris_cong_qcnn; henderson_2019_quanvo as an extension of the concept of classical NNs by utilizing quantum circuits as layers of the network. The design of the quantum layers corresponds to different ML architectures, with notable examples including Quantum Convolutional NN (QCNN) iris_cong_qcnn and Quanvolutional NN (QuanNN) henderson_2019_quanvo(see also recent advancements in QuanNN Bhatia_2023; Ceschini2025). While convolutional NNs (CNNs) – the cornerstone of modern NNs – have achieved significant success across a wide range of ML applications cnns_survey; cnns_survey2024, their resource requirements, especially for training large models, increase drastically with the size and complexity of data comp_limits_dl; dl_scaling_2017. Quantum NN extensions, such as QCNN and QuanNN, were proposed under the hypothesis that, since quantum computers can process information in fundamentally different ways, quantum NNs may potentially offer greater efficiency compared to their classical counterparts sup_learning_with_QC; ai_in_quantum.
@@
In parallel to developing quantum NN models, in recent years there has been a growing interest in creating a hybrid quantum-classical model where quantum and classical NNs work in tandem hybrid_qc_compare; hqcnn; hqcnn_downlink; hqcnn_vqc. This division of computing resources allows leveraging the strengths of both classical and quantum computing while managing the limitations of near-term quantum devices. Hybrid models can be particularly promising in domains such as image processing and generative modeling, where quantum circuits may offer advantages in high-dimensional feature spaces, while classical systems handle more routine tasks. Thus, the hybrid quantum-classical ML approach provides a feasible pathway toward enhancing ML tasks under the current quantum technological constraints.
@@
A typical hybrid model includes classical layers for pre-processing data, followed by quantum processing layers (circuits), and concludes with classical post-processing steps. Nevertheless, various strategies have been proposed for integrating quantum layers into classical deep NN architectures. For example, in image classification tasks, quantum layers are often positioned near the end of the hybrid network to generate the final output label Senokosov_2024, while some approaches involve encoding data through quantum layers at the beginning of the network img_class_hqcnn. The design of the quantum layer itself varies widely as well, with notable examples including QCNN iris_cong_qcnn and QuanNN henderson_2019_quanvo. These variations underscore the need for a comprehensive study of optimal quantum layer placement in hybrid networks, as well as a systematic approach to constructing effective quantum layers.
@@
In this work, we take a step forward in these two directions. We introduce a novel quantum learning architecture, which we term quantum feature extraction, or QuFeX for short, which combines and leverages techniques from QCNN iris_cong_qcnn and QuanNN henderson_2019_quanvo. At a high-level it integrates the data analysis structure of QCNN, which was shown to excel in classification tasks elhag2024quantumconvolutionalneuralnetworks; qnn_2018; Hur2021QuantumCN, with the data pipeline of QuanNN which is well suited for local transformation of the input data henderson_2019_quanvo. This combination allows us to design a quantum layer that can extract features from the data in a reduced dimensional representation of it. This in turn results in a reduced number of parallel quantum circuit evaluations compared to QuanNN. Thus, one of the main advantages of the QuFeX design is in its ability to be integrated in a hybrid fashion with classical layers, which makes it an excellent candidate for integration in deep classical NNs.
@@
As a proof of concept, we propose a novel hybrid quantum-classical model, a Qu-Net, were the QuFeX is integrated at the bottleneck of a classical U-Net model. U-Nets are a type of convolutional neural network architecture primarily designed for image segmentation tasks unet. They are widely used in applications such as medical imaging unet; Yuan2024, autonomous driving unet_carvana, and satellite image analysis unet_satellite, where precise pixel-level predictions are essential. We tested the performance of the proposed Qu-Net on three image segmentation datasets, FruitSeg30 fruitseg30, PH2 ph2, and ISBI-2012 EM isbi and compared it to that of an all-classical U-Net baseline. Our findings, reported in detail below, provide numerical evidence that with a careful design of quantum learning circuits, and with a strategic placement of those within deep NNs we can achieve quantum-classical hybrid learning models that leverage the strengths of both computing paradigms. Our demonstrations represent one of the first applications of quantum-enhanced deep NNs to real-world image segmentation tasks.
@@
The remainder of the paper is organized as follows. In Section II, we first present a brief overview of CNNs and their quantum counterpart architectures, setting the stage for the introduction of the QuFeX module. In Section III, we introduce the Qu-Net architecture utilizing the QuFeX at the bottleneck of the network, and in Sec. IV we numerically test and analyze the performance of the Qu-Net for various model sizes and datasets and compare it to a U-Net baseline. Finally, we offer conclusions and outlook in Section V.
@@ II.1 Convolutional neural networks and their quantum counterparts
Motivated by these powerful information processing components researchers have developed quantum counterparts of CNN. QCNN, first introduced in iris_cong_qcnn, was designed to mimic the translational invariance property of CNN. Figure 2 provides an illustration of the QCNN architecture. In QCNNs, the (classical) convolutional layers are replaced with a set of parametrized unitary transformations ($U_{j}$ in the figure) and pooling layers are replaced with fixed unitary operations ($V_{j}$ in the figure) conditioned on the measurement results of few specific qubits. These blocks of unitary operations (circuits) are concatenated in a serial fashion, and as shown in Fig. 2, each unitary circuit is designed in a translational invariant way to maintain this CNN property. One of the advantages of this approach is that for an $N$ qubit circuit, this network has only $O(\log N)$ trainable parameters iris_cong_qcnn. While the structure of each circuit layer is fixed, the parametrized unitary operations ($U_{j}$) themselves are trainable via an optimization strategy, e.g., gradient descent to minimize a particular loss function, and therefore are often thought of as the quantum-equivalent of CNN kernels.
While, operationally, the quantum convolutional circuits within QCNNs are considered efficient data processing modules that mimic the translational invariance of kernels in CNN, the data pipeline of QCNN generally does not possess some of the key features of CNNs. Specifically, since the convolutional and pooling layers in a typical QCNN circuit are ordered sequentially a circuit, the input to the circuit is often too large to be handled efficiently (this is known also as the input problem). On the other hand, classically manipulating the input to reduce it to a manageable size, may result in loss of information when handling data with multiple (or non-local) features.
@@
To address these challenges, QuanNN henderson_2019_quanvo has been proposed, which leverage quantum circuits to enable the simultaneous extraction of local features across input data. QuanNNs act as feature extractors, analogous to convolutional layers in classical networks, and can be integrated seamlessly with classical NN architectures for hybrid quantum-classical processing. As illustrated in Fig. 3, in QuanNN each quantum circuit is treated as a kernel and generates a feature map when applied to an input by transforming spatially-local subsections of it. These filters correspond to specific circuit designs and can be configured to potentially learn different features of the input. The QuanNN architecture thus offers greater flexibility and is considered a closer representation of classical CNNs in its way of handling the flow of data. Nonetheless, by mimicking the data flow of classical CNN, QuanNN fall short in efficiency. Since QuanNNs extract local quantum features independently for each patch, similar to CNNs, to process the input data we must design ${\cal O}(kN)$ quantum experiments, with thousands of learning iteration per filter as reported in henderson_2019_quanvo, where $N$ is the size of the input and $k$ is the number of filters used.
@@ II.2 QuFeX: Quantum feature extraction architecture
In this architecture, the QuFeX circuit is designed as a quantum filter that creates a feature map from its input data. We note by passing that while tensor-network architectures can also act as efficient classical filters through tensor decomposition, QuFeX differs conceptually and operationally: it implements a trainable quantum circuit whose filtering behavior arises from coherent operations on entangled states. Rather than claiming superiority over classical tensor-network schemes, we emphasize QuFeX as a proof-of-concept demonstration showing that shallow quantum circuits, even with few qubits, can enhance feature extraction within established deep-learning frameworks.
[…]
Figure 4: Example of a QuFeX architecture. The QuFeX incorporates ideas and techniques from QCNN and QuanNN. In the data flow we use information taken from different feature maps (from two maps in this example) and pass it to a convolutional-inspired quantum circuit. Example of such circuit is illustrated on the top-right part of the figure where $U_{1}$ and $U_{2}$ depend on trainable parameters (denoted in the paper by $\vec{\theta}$), see Appendix A for more details. Single-qubit measurements are then performed and the average of the corresponding Pauli-$Z$ operator is then passed on to produce an output feature map. In the QuFeX architecture we use different circuits (i.e., circuits on different qubits) to analyze different patches of the input data, in parallel. The circuits that run in parallel contain the same trainable set of parameters.
Within each QuFeX layer we process the data in a translationally invariant way, similar to QCNNs, as depicted in the Fig. 4, with one critical change. Here, we are not discarding the control qubits in the pooling layer. In QCNN the pooling layers are implemented by discarding the measured (control) qubits, effectively reducing the dimension of the network. In the QuFeX architecture, similar to classical CNNs and QuanNN, the reduction of dimension is done directly through the data flow. Therefore, instead of discarding the control qubits, and by doing so discarding information within a filter, we keep the control qubits and treat their output as a part of the (or even separate) feature map.
@@
As a proof of concept, in the next sections we propose a novel hybrid quantum-classical model, a Qu-Net, were the QuFeX is integrated within a classical U-Net model unet— a CNN with a U-shape topology specifically developed for image segmentation. We demonstrate the potential power of the QuFeX though segmentation performance of the Qu-Net architecture. Our demonstrations represent one of the first applications of quantum-enhanced deep NNs to real-world image segmentation tasks.
III The Qu-Net architecture
@@ III.1 The U-Net architecture
Illustrated in Fig. 5(a), the U-Net architecture is a deep CNN with a U-shape topology specifically developed for image segmentation unet. It features an encoder-decoder structure with skip connections, which help preserve spatial information, and enable precise pixel-level predictions for more accurate segmentation results. The encoder progressively downsamples the input image, capturing essential features through convolutional and pooling layers. Subsequently, the decoder upsamples the compressed feature map to the original input size, reconstructing the segmentation map. At its bottleneck, i.e., at the most compressed layers, the U-Net transforms features into a latent representation that bridges the encoder and decoder. This design “encourages” the model to retain only the relevant information in the bottleneck layer, facilitating the construction of an effective segmentation map.
This design helps U-Net excel in segmenting objects with fine boundaries, making it highly effective, for example, in medical applications like tumor delineation, organ segmentation, and cell tracking unet; Yuan2024. Beyond medical imaging, U-Net is used in various fields requiring pixel-level classification, e.g., for satellite image analysis unet_satellite and autonomous driving unet_carvana. Its ability to handle small datasets through data augmentation and its versatile performance in capturing detailed structural information has made U-Net a widely adopted tool in image segmentation across industries.
@@ III.2 The Hybrid Qu-Net architecture
The downscaling of the input all the way down to the bottleneck of the architecture, introduces limitations on the representability of the original image. In addition to QuFeX, inspired by the ResNet architecture resnet, we include (classical) residual connections from one part of the network to the other that bypass the QuFeX module, see Fig. 5(b). The residual connection can be expressed as $\mathbf{y}=\mathcal{Q}(\mathbf{x})+\mathbf{x}$ where $\mathcal{Q}(\mathbf{x})$ represents the output of the quantum layer applied to the input $\mathbf{x}$ and the addition of $\mathbf{x}$ serves as the identity mapping. We argue that residual connections, typically used to stabilize training and convergence in deep NNs, are very relevant in hybrid quantum-classical setups. Adding identity mapping over the quantum layer enables the model to propagate features learned by the classical layers forward and backward through the network without being disrupted by the quantum transformation. This configuration allows the network to select an optimal path: either through a classical-only route or a path incorporating quantum processing, thus leveraging the advantages of both. Our tests below show that introducing residual connections to the Qu-Net architecture provides better segmentation performance compared to baseline U-Net models where the QuFeX is not included.
IV Testing and evaluation
IV.1 FruitSeg30 Image Segmentation
@@
We have tested and evaluated the performance of the proposed QuFeX module within the novel Qu-Net architecture using the FruitSeg30 Segmentation Dataset & Mask Annotations fruitseg30 which consists of high-resolution images of a variety of fruits and their corresponding segmentation masks. The complete dataset includes 30 fruit classes, totaling 1969 images, each paired with precise segmentation masks. For this research, a curated subset of $751$ images was selected (along with their segmentation masks) and downscaled, using Bilinear interpolation, from the original $512\times 512$ pixels to $64\times 64$ pixels to align with the available computational resources [Intel Core i5-11300H CPU @ 3.10GHz (4 cores), 8 GB DDR4 RAM running Ubuntu 22.04 LTS]. The subset was selected to ensure diversity between fruit types while maintaining computational feasibility.
To provide a baseline for the performance of the Qu-Net architecture, a U-Net architecture with approximately the same number of trainable parameters was also trained and evaluated on the same dataset. The code for all the numerical experiments reported in this subsection are available in a public repository Naman2025.
We tested three variations of the models, which we refer to as tiny, small, and medium, that consist of approximately 12,000, 26,000, and 40,000 trainable parameters, respectively. These variations in parameter count were achieved by adjusting the number of filters in the classical layers of the architectures. We refer the reader to the Appendix A for more details about the specific U-Net and Qu-Net models design.
@@
Our models were trained using an Adam optimizer with a learning rate of $0.001$ and a binary cross-entropy loss function. Training was performed for 10 epochs with a batch size of 64. The Qu-Net architecture was executed on quantum simulator using PennyLane pennylane. An ablation study was conducted to evaluate the influence of the QuFeX layer’s design choices, see the appendix A for more details. The performance of Qu-Net was analyzed with varying numbers of quantum filters and qubits per filter within the QuFeX module. Specifically, here we report the results obtained using one quantum filter with 8 qubits and 2 quantum filters with 4 qubits per filter. We denote those models by Qu-Net $8(1)$ and Qu-Net $4(2)$, respectively.
@@
The statistical summary of these metrics is visualized in Figs. 7 and 8, providing a comparative analysis of the models’ performance. Each boxplot depicts the distribution of the results (IoU) from the 10 independent runs for the tested models. The median value is marked, with the interquartile range (IQR) capturing the variability of performance across different dataset partitions. Outliers are displayed to identify cases where the model’s performance deviates significantly from the norm. The complete distribution of results is given in Fig. 15 in the appendix A.